Physics Midterm Free Fall and Universal Gravitation 19. Which of the graphs above accurately display a distance- time graph for an object in freefall, assume the graph starts at the point that the object was dropped, and that downward motion is positive? 20. Which of the graphs above accurately display a Velocity -time graph for an object in freefalil, assume the graph starts at the point that the object was dropped, and that downward motion is positive? 21. Which of the graphs above accurately display an acceleration - time graph for an object in freefall, assume the graph starts at the point that the object was dropped, and that downward motion is positive? planets of equal radius have different accelerations due to gravity (g). Planet As g 10 m/s, Planet Bs g is 20 much more massive is Planet B compared to Planet A? A. Planet A is 2 times as massive as Planet B B. Planet A has the same mass as Planet B C. Planet B is 2 times as massive as Planet A D. Not enough information provided 23. Two objects of different mass are dropped in a vacuum. Which of the foilowing statements is true? A. All objects will accelerate at the same rate downward. B. All objects will accelerate at the different rates dowaward. C. Heavier objects fall faster than light objects D. lighter objects fall faster than heavy objects 24. What is the correct relationship between gravitational force and the distance between these two masses? A. inversely proportional to distance squared B. directly proportional to distance C. inversely proportional to mass squared. D. directly proportional to the sum of the square root
Answer #1

The distance time graph will be the graph (a) , the velocity vs time graph will be graph (c) as object starts from rest and velocity increase, acceleration vs time graph will be graph (b) as acceleration is constant.

--------------------------------------------------------------------------------------

we know that

g = GM /r2

as radius is same, we can write

g = M (G is a constant)

M = gr2 / G

planet B is 2 times more massive than planet A

------------------------------------------------------------------------------

In vacuum, all objects accelerate at same rate

-----------------------------------------------------------------------------

F = Gm1m1 / r2

inversely proportional to distance squared.